the probability of getting the sum of 5 when you roll two number cubes is 4/36. how many times should you expect to get a sum of 5 if you roll the cubes each number of times? a.9 b.18 c.27 d.100 d.450
Answer by edjones(8007)

a. 9* 4/36 = 36/36 = 1 time
b.18 * 4/36 = 72/36 = 2 times
c. 27=3 times
d. 100 * 4/36 = 400/36 = 11.11 times
e. 450 * 4/36 = 1000/36 = 27.78 times
.
Ed